OpenStudy (anonymous):

A ball moving at +3.0 m/s along a table rolls off a table and lands on the ground 2.0 m away. How high was the table?

OpenStudy (shamim):

Can u draw it

OpenStudy (shamim):

U know there is no x component of gravitational acceleration g

OpenStudy (shamim):

So x=V0*t 2=3*t t=?

OpenStudy (anonymous):